# Decentralization of Ethereum’s Builder Market

### v1

@@ 6. Related Work
Empirical study on Ethereum’s builder market.   Several previous works studied the centralization of the builder market shares. An early exploration by Yang et al. (Yang et al., 2022) analyzed the market share of relays and builders. Their findings indicated a centralization of Ethereum’s builder market, with Flashbots and builder0x69 accounting for over 53% of the market share between September 15th, 2022, and November 30th, 2022. Similar studies on the builder market conducted by (Wahrstätter et al., 2023b; Heimbach et al., 2023) also confirmed the centralization in the builder market. Their observations were based on the market share and the distribution of total profits generated within the PBS ecosystem. Moreover, several online dashboards (Wahrstätter, 2024; Labs, 2024; Bitfly, 2024; Flashbots, 2024) offer a dynamic view of the ecosystem’s trends by presenting real-time PBS analytics. However, these studies do not quantify the causes of centralization or the security implications of a centralized builder market.
@@
Private order flow.   Many studies also discuss the importance of private order flows, particularly focusing on private order flows from integrated searchers. Thiery (Thiery, 2023) counted the total number of transactions, along with their values, from June 1 to July 15, 2023. He found that private order flows account for only 30% of the transactions yet represent 80% of the total value paid to builders. Our study highlights a similar result: in over 50% of blocks, private order flows account for more than 60% of the block value. Gupta et al. (Gupta et al., 2023) confirmed that builders are more likely to win the MEV auction when their integrated searchers can provide high-value exclusive order flow by effectively exploiting arbitrage between a centralized exchange (CEX) and a decentralized exchange (DEX). The dashboard by Winnie (Xiao, 2024) and the study by Titan (Titan, 2023) analyzed the vertical integration between searchers and builders through the quantity and value of transactions sent from searchers to builders. Heimbach et al. (Heimbach et al., 2024) observed that during their study, Beaver Builder received a total of 1,941.1 ETH in transaction fees but paid proposers 6,620.94 ETH. This large gap suggests that Beaver Builder might have received significant off-chain profits from its integrated searchers. These studies show that three integrated builders (Beaver, Rsync, and Manta) gain a significant advantage in order flow from their integrated searchers. Our study made the same observation and further quantified the pivotal level.
@@
Strategic behaviors in MEV-Boost auctions.   The strategies used by different entities in the MEV-Boost auctions and their potential impacts have attracted significant interest from researchers. The strategic behaviors of proposers were first studied by (Schwarz-Schilling et al., 2023; Wahrstätter et al., 2023b) focusing on the timing game where proposers strategically delay block proposals to maximize their profits. Öz et al. (Öz et al., 2023) further conducted an agent-based simulation to analyze how waiting games affect consensus stability, finding that a delay strategy can be profitable and not degrade consensus if sufficiently many validators adopt it. A recent study (Always, 2024) empirically investigated timing games in the wild and how block proposal delays affect the block fork rate. The results indicate that three validator sets–P2P, Kiln, and Attestant, that are known or suspected of playing timing games–did not delay their block proposals enough to cause their blocks to be forked. Wu et al. (Wu et al., 2024) proposed a game-theoretic model for MEV-Boost auctions, using simulations to examine various builders’ bidding strategies, including naive, adaptive, last-minute, and bluff bidding. Their results highlighted the importance of latency and exclusive order flow in the effectiveness of bidding strategies. Pai et al. (Pai and Resnick, 2023) analyzed the latency advantage of searcher-builder integration in adjusting bidding values. The strategies discussed in these papers highlight the potential factors that can make MEV-Boost auctions uncompetitive or inefficient, while our paper quantifies these uncompetitive and inefficient auctions.
@@
Causes and implications of centralized builder market.   A provider may exclusively share its order flow with a single builder. Previous works such as (Kilbourn, 2022; Frontier, 2023; Gupta et al., 2023) pointed out that exclusive order flows have potential centralizing effects on the builder market, because builders with exclusive order flows are more likely to win MEV-Boost auctions, thus gaining further advantages in receiving order flows from non-integrated searchers and in winning order flow auctions. The decentralization of Ethereum’s builder market not only benefits itself but also positively impacts the decentralization of validators. A recent study showed that heterogeneity among validators in building blocks may lead to centralization; however, a decentralized and competitive builder market can encourage validator decentralization (Bahrani et al., 2024). We complement these theoretical analyses with empirical studies to analyze potential problems caused by the current centralized builder market.
@@
Blockchain decentralization metrics.   Nakamoto coefficient was proposed to describe the minimum number of entities required to disrupt the blockchain’s subsystem (Srinivasan and Lee, 2017). Other metrics, such as the Gini coefficient (Dorfman, 1979), Shannon entropy (Lin, 1991), and Herfindahl–Hirschman index (Rhoades, 1993), have also been utilized to measure decentralization (or the lack of it) in staking distribution (Grandjean et al., 2023; Kwon et al., 2019), mining power (Lin et al., 2021; Wu et al., 2019; Kwon et al., 2019), and block building (Heimbach et al., 2023). An SoK paper (Zhang et al., 2022) comprehensively summarized the existing works on blockchain centralization; however, none explored decentralization from the perspective of entry barriers or auctions.
7. Discussion
7.1. Limitations of Our Approach and Data
@@
True value.   We compute the true values of bids using a similar approach as that in MEV analysis tools (EigenPhi, 2023; Flashbots, 2023b) and previous works (Heimbach et al., 2023; Thiery, 2023). This approach has a few limitations. First, it relies on the knowledge of the set of addresses controlled by each builder. Most builders use well-known addresses to receive payments from searchers and make payments to proposers, but it is possible that some builders use less publicized addresses. We cross-validated our builder address dataset with other sources to ensure completeness (see Sec. 3.1), but if we missed a builder’s address, the computation of extracted value (and hence true value) would be inaccurate. Secondly, this approach neglects off-chain profits, such as off-chain payments received from searchers. Off-chain payments are more likely to happen between integrated builder-searcher pairs. We might underestimate the true value of a block due to unknown off-chain payments, which could lead to the misclassification of an efficient auction as inefficient due to subsidization. However, in our study, we did not observe any excessively large subsidies. Furthermore, underestimating the winner’s true value does not affect our assessment of competitiveness. We might misclassify an uncompetitive auction as competitive if we underestimate the second-highest true value. However, such underestimation should be uncommon, as we observe that most builders bid around our computed true value. It would be irrational for builders not to bid higher to win the auction if they actually have a higher true value.
@@
Missing pivotal providers.   Our results in Sec. 4.1 might omit some pivotal providers since the providers we can identify are a subset of all providers. For example, BloXroute (a widely used channel) does not publish the list of transactions they processed. Thus, we can’t evaluate its pivotal level. This limitation is common in existing studies (Lu, 2024; Lu, Angela and Sui, Danning and Durnford, Jaden, 2024). However, there will not be false positives (classification of a non-pivotal provider as pivotal) because we identify a given transaction’s provider based either on the official dataset from MEV-Share and MEV Blocker or on the specific searchers’ addresses (see Sec. 3). Thus, the significance of each provider will only be underestimated rather than overestimated.
@@
7.2. Future Works
Based on the findings, we now outline directions to enhance the decentralization of builder markets. These directions present interesting technical challenges that we leave for future work.
Several issues we identified boil down to the lack of a trustworthy mechanism to facilitate fair exchanges between private order flow providers and builders. In today’s ecosystem builders are at an advantage. To protect themselves, private order flow providers resort to reputation-based solutions such as requiring a minimal market share, which creates siloes.
One approach is building platforms to facilitate the said fair exchange. For example, MEV-Share is one such platform between searchers and builders. To receive order flows from MEV-Share, builders must agree to the Fair Market Principles set force by the community (Flashbots, 2023a). Should builders fail to adhere to these principles, they will be either removed from or flagged by the community.
However, in its current form, MEV-Share is a trusted party. One direction being explored (e.g., by SUAVE (Flashbots, 2024c)) is to use Trusted Execution Environments (TEEs) to run block-building algorithms without leaking private order flows to the builders, but standard challenges such as side channels and covert channels (Salles et al., 2023) apply.
Another unsolved problem is how to reliably detect private order flow leakage across builders. Currently, after receiving private orders, a builder can leak the orders to another builder (potentially a Sybil), bypassing any reputation or economic penalty. One idea is to embed watermarks in order flows to enable tracking. Watermarks must be invisible to the adversary and not interfere with transaction execution. Designing a secure watermarking scheme against a practical adversary model is an interesting future direction.
A reliable leakage detection mechanism (e.g., watermarking) will open up the design space of reputation systems. For example, one possibility is to ask builders to deposit a certain amount of collateral to access private order flows, and order flow leakage will be penalized economically.
8. Conclusion
In this paper, we characterized the efficacy of PBS by evaluating its openness to new builders and assessing the competitiveness and efficiency of the auctions. Our study identifies a significant entry barrier in the builder market: private order flow. We further quantify the barrier costs—a new builder must pay up to 1.4 ETH to access this flow. Given this barrier, it is not surprising that we observed varying capabilities among builders in extracting MEV, with inequalities among top, middle, and tail builders worsening as MEV increases. Our study shows that 88% of the MEV-Boost auctions were competitive, while 79% were efficient. The proposers’ losses due to uncompetitive auctions are 0.98% of their total gain. Based on these insights, we propose suggestions for future developments to mitigate the trust crisis in the MEV-Boost auctions and facilitate the decentralization of Ethereum’s builder market.
@@ Acknowledgements
This work is partially supported by an Ethereum Foundation Grant. We thank ultra sound relay for sharing the full bid dataset. We also thank the teams of BloXroute, Merkle, MeowRPC, MEV Blocker, Flashbots, and Blocknative for answering our questions about access requirements. We are grateful for the valuable discussions with Thomas Thiery and Barnabé Monnot.
@@ Appendix A Visualization for distribution
Figure 10. Distribution of the percentage of builders who submit to the ultra sound relay.
Figure 11. CDF of the winning bid value of auction dataset.
Appendix B List of Searchers and Builders
In Tab. 4, we provide the list of searchers and their addresses mentioned in our study. In Tab. 5, we have counted the top 25 builders in the market, ranked by their market share from April 2023 to August 2023.
Table 4. The list of searchers
@@ Appendix B List of Searchers and Builders
Table 5. The percentages of blocks and MEV-Boost blocks built by each builder from April to August 2023.
@@
Beaver 22.418739 24.719523
Rsync 18.611899 20.521996
Builder0x69 17.799124 19.625807
Flashbots 13.276622 14.639172
Titan 7.105763 7.835011
BloXroute 2.384068 2.628740
Blocknative 2.267458 2.500162
F1b 2.158849 2.380407
eth-builder 1.086828 1.198366
BuildAI 0.894899 0.986741
payload.de 0.645310 0.711536
lightspeedbuilder 0.321781 0.354805
boba-builder 0.257682 0.284128
Manta 0.254004 0.280072
Eden 0.202596 0.223388
0x3be…436 0.186227 0.205339
It’s Free Real Estate 0.133072 0.146728
Gambit Labs 0.117806 0.129896
bobTheBuilder 0.102815 0.113367
Manifold 0.088929 0.098056
blockbeelder 0.084239 0.092884
finest artisanal blocks 0.074307 0.081933
nfactorial 0.072375 0.079803
0x563…cf3 0.041660 0.045935
uwu.builders 0.016370 0.018050
Appendix C Openness of the builder market
C.1. Importance of private order flows

### v3

@@ 6 Related Work
Empirical study on Ethereum’s builder market.   Several previous works studied the centralization of the builder market shares. An early exploration by Yang et al. [92] analyzed the market share of relays and builders. Their findings indicated a centralization of Ethereum’s builder market, with Flashbots and builder0x69 accounting for over 53% of the market share between September 15th, 2022, and November 30th, 2022. Similar studies on the builder market conducted by [84, 41] also confirmed the centralization in the builder market. Their observations were based on the market share and the distribution of total profits generated within the PBS ecosystem. Moreover, several online dashboards [87, 49, 7, 36] offer a dynamic view of the ecosystem’s trends by presenting real-time PBS analytics. However, these studies do not quantify the causes of centralization or the security implications of a centralized builder market.
@@
Private order flow.   Many studies also discuss the importance of private order flows, particularly focusing on private order flows from integrated searchers. Thiery [80] counted the total number of transactions, along with their values, from June 1 to July 15, 2023. He found that private order flows account for only 30% of the transactions yet represent 80% of the total value paid to builders. Our study highlights a similar result: in over 50% of blocks, private order flows account for more than 60% of the block value. Gupta et al. [40] confirmed that builders are more likely to win the MEV auction when their integrated searchers can provide high-value exclusive order flow by effectively exploiting arbitrage between a centralized exchange (CEX) and a decentralized exchange (DEX). The dashboard by Winnie [91] and the study by Titan [81] analyzed the vertical integration between searchers and builders through the quantity and value of transactions sent from searchers to builders. Heimbach et al. [42] observed that during their study, Beaver Builder received a total of 1,941.1 ETH in transaction fees but paid proposers 6,620.94 ETH. This large gap suggests that Beaver might have received significant off-chain profits from its integrated searchers. These studies show that three integrated builders (Beaver, Rsync, and Manta) gain a significant advantage in order flow from their integrated searchers. Our study made the same observation and further quantified the pivotal level.
@@
Strategic behaviors in MEV-Boost auctions.   The strategies used by different entities in the MEV-Boost auctions and their potential impacts have attracted significant interest from researchers. The strategic behaviors of proposers were first studied by [76, 84] focusing on the timing game where proposers strategically delay block proposals to maximize their profits. Öz et al. [66] further conducted an agent-based simulation to analyze how waiting games affect consensus stability, finding that a delay strategy can be profitable and not degrade consensus if sufficiently many validators adopt it. A recent study [1] empirically investigated timing games in the wild and how block proposal delays affect the block fork rate. The results indicate that three validator sets–P2P, Kiln, and Attestant, that are known or suspected of playing timing games–did not delay their block proposals enough to cause their blocks to be forked. Wu et al. [90] proposed a game-theoretic model for MEV-Boost auctions, using simulations to examine various builders’ bidding strategies, including naive, adaptive, last-minute, and bluff bidding. Their results highlighted the importance of latency and exclusive order flow in the effectiveness of bidding strategies. Pai et al. [67] analyzed the latency advantage of searcher-builder integration in adjusting bidding values. The strategies discussed in these papers highlight the potential factors that can make MEV-Boost auctions uncompetitive or inefficient, while our paper quantifies these uncompetitive and inefficient auctions.
@@
Causes and implications of centralized builder market.   A provider may exclusively share its order flow with a single builder. Previous works such as [46, 38, 40] pointed out that exclusive order flows have potential centralizing effects on the builder market, because builders with exclusive order flows are more likely to win MEV-Boost auctions, thus gaining further advantages in receiving order flows from non-integrated searchers and in winning order flow auctions. The decentralization of Ethereum’s builder market not only benefits itself but also positively impacts the decentralization of validators. A recent study showed that heterogeneity among validators in building blocks may lead to centralization; however, a decentralized and competitive builder market can encourage validator decentralization [4]. We complement these theoretical analyses with empirical studies to analyze potential problems caused by the current centralized builder market.
@@
Blockchain decentralization metrics.   Nakamoto coefficient was proposed to describe the minimum number of entities required to disrupt the blockchain’s subsystem [77]. Other metrics, such as the Gini coefficient [18], Shannon entropy [51], and Herfindahl–Hirschman index [73], have also been utilized to measure decentralization (or the lack of it) in staking distribution [39, 47], mining power [52, 89, 47], and block building [41]. An SoK paper [93] comprehensively summarized the existing works on blockchain centralization; however, none explored decentralization from the perspective of entry barriers or auctions.
7 Discussion
7.1 Mitigation
Our results suggest that proposers suffer from significant losses because of builder centralization, which can lead to two undesirable outcomes: First, PBS could become unstable. Second, MEV measured by the auctions may be inaccurate.
In this section, we propose a structural change to the existing MEV supply chain Sec. 2.2 and a solution based on the new supply chain structure to address these two problems.
In the new supply chain, only one builder receives order flows from providers, and proposers can only accept blocks from this builder.
Protocol.   Our solution involves having the builder run the truthful mechanism in [57] in a Trusted Execution Environment (TEE) [15] The pseudocode for the TEE builderis described in Fig. 17.
The builder provides two functions. First, it opens RPC endpoints to receive the private order flows from the providers. Providers send their order flows along with a bid indicating the MEV in their order flows. Second, when it receives a signal for block building (e.g., a new slot starts in Ethereum), it will run an ordering algorithm to build a block with optimal MEV based on the received order flows (a reasonably performant ordering algorithm is described in Appx. E).
The TEE builder runs the truthful mechanism in [57]: it refunds all providers for their marginal value contribution to the block. For each included order flow, the builder builds an alternative block without that specific order flow to calculate its marginal value contribution by comparing the MEV extracted between these two blocks. Next, we outline the key advantages of this solution.
Security guarantees.   TEEs offer a secure and isolated environment, ensuring that private order flows remain confidential and protected from external tampering. Moreover, order flow providers can utilize remote attestation to verify that the builder running within the TEE is indeed the intended one, thereby effectively mitigating trust concerns.
Incentive compatibility.   [57] shows that for providers, truthfully bidding is a dominant strategy when they are refunded based on their marginal value contribution. The formal proof and the required assumptions are in [57]. Here, we offer an intuitive explanation: a provider effectively pays the second price for their order flows when refunding their marginal value contribution.
Analysis.   Since truthfully bidding is an incentive-compatible dominant strategy for all providers, the TEE builder can compute the total MEV in a block from the providers’ bids, achieving an accurate MEV oracle. Moreover, because the builder program in the TEE can be programmed to pay the proposer fairly, the instability of PBS is also addressed.
Future works.   We leave the analysis of whether the new supply chain structure is sustainable in the long term for future work. Intuitively, since providers are incentivized to send order flows only to the builder due to refunds, other builders will naturally be driven out of the market, leaving proposers to only accept blocks from the TEE builder. Furthermore, since real-world proposers might behave rationally, exploring the new equilibrium that emerges among a rational provider, builder, and proposer is another intriguing question to be addressed. Finally, this solution provides an accurate MEV oracle and transparent MEV information. These insights open up new possibilities for future MEV mitigation directions.
7.2 Limitations of Our Approach and Data
@@
True value.   We compute the true values of bids using a similar approach as that in MEV analysis tools [20, 31] and previous works [41, 80]. This approach has a few limitations. First, it relies on the knowledge of the set of addresses controlled by each builder. Most builders use well-known addresses to receive payments from searchers and make payments to proposers, but it is possible that some builders use less publicized addresses. We cross-validated our builder address dataset with other sources to ensure completeness (see Sec. 3.1), but if we missed a builder’s address, the computation of extracted value (and hence true value) would be inaccurate. Secondly, this approach neglects off-chain profits, such as off-chain payments received from searchers. Off-chain payments are more likely to happen between integrated builder-searcher pairs. We might underestimate the true value of a block due to unknown off-chain payments, which could lead to the misclassification of an efficient auction as inefficient due to subsidization. However, in our study, we did not observe any excessively large subsidies. Furthermore, underestimating the winner’s true value does not affect our assessment of competitiveness. We might misclassify an uncompetitive auction as competitive if we underestimate the second-highest true value. However, such underestimation should be uncommon, as we observe that most builders bid around our computed true value. It would be irrational for builders not to bid higher to win the auction if they actually have a higher true value.
@@
Missing pivotal providers.   Our results in Sec. 5.1 might omit some pivotal providers since the providers we can identify are a subset of all providers. For example, BloXroute (a widely used channel) does not publish the list of transactions they processed. Thus, we can’t evaluate its pivotal level. This limitation is common in existing studies [53, 54]. However, there will not be false positives (classification of a non-pivotal provider as pivotal) because we identify a given transaction’s provider based either on the official dataset from MEV-Share and MEV Blocker or on the specific searchers’ addresses (see Sec. 3). Thus, the significance of each provider will only be underestimated rather than overestimated.
@@
7.3 Security Problems in PBS
As we mentioned in Sec. 5.2, one reason for the access barrier to private order flows is the difficulty of establishing trust. A malicious builder, relay, or proposer may attack providers. A real-world incident involved a malicious proposer who double-signed a header to unbundle a builder block and further stole MEV from searchers by proposing a different block [61].
Another security concern is censorship. Ethereum’s censorship resistance currently depends heavily on whether builders choose to censor, particularly the actions of the top builders. For example, before the top builder, Beaver, stopped censoring in July 2024, over 50% of blocks were censoring [86]. Our study reveals that builder centralization remains difficult to overcome, potentially exacerbating the censorship problem in Ethereum.
8 Conclusions
In this paper, we interrogate the causes and implications of builder centralization in Ethereum and challenge the community’s belief that it is acceptable. Our empirical study shows that the current blocker market has led to significant profit losses for validators. These significant losses could lead to PBS’s instability and MEV oracles’ inaccuracy and, if left uncontrolled, can lead to validator centralization. We attribute the loss to two factors: insufficient competition and inequality of block-building capacities, with the latter being the dominant reason. We further investigated the trust and incentive issues of the current MEV supply chain that contribute to builder centralization. To address these two problems, we propose a structural change to the current MEV supply chain and a solution based on this new structure.
@@ Acknowledgements
This work is partially supported by an Ethereum Foundation Grant. We thank ultra sound relay for sharing the full bid dataset. We also thank the teams of BloXroute, Merkle, MeowRPC, MEV Blocker, Flashbots, and Blocknative for answering our questions about access requirements. We are grateful for the insightful comments by Prof. Campbell Harvey and the valuable discussions with Thomas Thiery, Barnabé Monnot, Danning Sui, Burak Öz, Data Always, and Georgios Konstantopoulos.
@@ Appendix A Visualization for Distribution
Figure 10: Distribution of the percentage of builders who submit to the ultra sound relay.
Figure 11: CDF of the winning bid value of auction dataset.
Appendix B Data Collection
Builder public keys.   We collect builder public keys from two data sources. First, builders often disclose their public keys in official documents. For example, Titan builder has disclosed 15 public keys in its official document [10]. Second, some builders leave their unique marks in the blocks’ “extra data” field. For example, Flashbots builder uses the mark ‘‘Illuminate Democratize Distribute’’. Given that such marks are unauthenticated and susceptible to impersonation, we ignore them unless they are used more than 100 times in the bids by the same builder public key.
Transaction sources.   We identified transactions originated from Telegram trading bots, including Maestro [55], Banana Gun [5], and Unibot [82], by searching for transactions that interacted with their official smart contracts.
Two channels, MEV Blocker and MEV-Share publish data [70, 33] about transactions submitted to them. Specifically, transactions submitted to MEV Blocker are publicly available on Dune [70], from which we extracted 22,819,437 transaction hashes that fall within our study’s range. We obtained 20,334,513 transactions submitted to MEV-Share from the public Flashbots dataset [33]. These transaction labels are sourced from their official datasets and we view them as ground truth.
Appendix C List of Searchers and Builders
In Tab. 3, we provide the list of searchers and their addresses mentioned in our study. In Tab. 4, we have counted the top 25 builders in the market, ranked by their market share from April 2023 to December 2023.
Table 3: The list of searchers
@@ Appendix C List of Searchers and Builders
Table 4: The percentages of blocks and MEV-Boost blocks built by each builder from April to December 2023.
@@
Beaver 22.53 24.69
Rsync 20.75 22.74
builder0x69 12.65 13.86
Titan 12.35 13.53
Flashbots 12.24 13.42
BloXroute 2.14 2.34
f1b.io 1.66 1.81
Blocknative 1.40 1.53
eth-builder 0.77 0.84
BuildAI 0.73 0.80
Gambit Labs 0.64 0.70
Jetbldr 0.47 0.51
Manta 0.46 0.51
payload.de 0.39 0.42
boba-builder.com 0.27 0.30
Penguinbuild 0.23 0.25
Eden 0.22 0.24
lightspeedbuilder 0.21 0.23
0x3be…436 0.16 0.18
lokibuilder.xyz 0.12 0.13
bobTheBuilder.xyz 0.11 0.13
blockbeelder.com 0.11 0.12
finest artisanal blocks 0.09 0.10
It’s Free Real Estate 0.08 0.09
nfactorial.xyz 0.06 0.07
Appendix D Implications of Builder Centralization
D.1 Efficiency of MEV-Boost Auctions
Analysis of inefficiency.   Fig. 4 shows that 24.9% of MEV-Boost auctions were not efficient. In an inefficient auction, the builder with the highest true value does not win the auction. There are two possible reasons. First, a builder can win the auction without possessing the highest true value by overbidding, known as block subsidization [36]. Second, builders may “shade” their bids (bid shading means strategically placing bids slightly below one’s true value); if the builder with the highest true value “shades too much,” it will lose.
We analyzed the underlying reasons for all inefficient auctions. In 61.0% of cases, winners win the auctions due to subsidization. In 48.0% of cases, the builder with the highest true value shaded too much. It is important to note that auctions may exhibit inefficiency for multiple reasons. For instance, the winner may subsidize her bid while simultaneously, the builder with the highest true value shades. Thus, the cumulative percentage exceeds 100%.
It is important to note that inefficiency does not necessarily disadvantage the validators. In fact, validators may receive more revenues in inefficient scenarios involving subsidization.
MEV’s implication on efficiency.   The ratio of efficient auctions in each tier of MEV is shown in Fig. 5. We note that the trend of efficiency is less expected. First, only 56.6% of the auctions within the low MEV tier are efficient. This can be attributed to the generally low overall bid values, rendering the auction outcomes vulnerable to alterations from minor subsidies. Second, while auctions tend to be more efficient as MEV increases, we still observe that 77.1% of auctions in the high MEV tier are efficient, and about 86% of inefficient cases result from bid shading, where the builders with the highest true values place their bids much lower than their true values.
D.2 Measuring Block-building Capacities
Figure 12: Violin graphs for the distribution of the CV across three MEV tiers, for top, middle, and tail builders. The CV measures the inequality of block-build capacity across builders.
D.3 Implications of Uncompetitive Auctions
Figure 13: The distribution of the CIs of MEV-Boost auctions
Appendix E Reference Builder Algorithm
Flashbots has open-sourced a Rust implementation of the MEV-Boost builder [34], which serves as a reasonably performant reference for builder algorithms. We will briefly introduce their approach. The builder processes each bundle (where a single transaction is considered a bundle containing one transaction) by computing its effective gas price and total profit. Subsequently, the builder sorts all received bundles in descending order based on either effective gas price or total profit. Employing a greedy algorithm, the builder then appends bundles from this sorted list. 1 presents the pseudocode of their algorithm.
Listing 1: Flashbots block-building algorithm
# Compute the effective gas price and profit for a given bundle
# A single transaction is recognized as a bundle of size 1
def receive(bundle):
effective_gas_price, profit = compute_price_and_profit(bundle)
bundle.egp = effective_gas_price
bundle.profit = profit
return bundle
# Building a block
def build(bundles, sorting_key):
# Sort bundles by price or profit
if sorting_key == ’egp’:
bundles.sort(key=lambda b: b.egp)
else:
bundles.sort(key=lambda b: b.profit)
# Initialize
B = []
# Get the current state
state = get_current_state()
# Greedy algorithm
while not bundles.is_empty():
# Get the first bundle from the queue
bundle = bundles.pop(0)
# Simulate the execution of the bundle
success = simulate_bundle(state, bundle)
# Update with the successful bundle
if success:
B.append(bundle)
state.update(bundle)
# Compute payment based on the final state
payment = compute_payment(state)
# Append the payment to the block
B.append(payment)
# Return the valid block
return B
Appendix F What Caused Inequality in Block-building Capacities
F.1 Importance of Private Order Flows